### v1

# Convergence of the extended Kalman filter with small and state-dependent noise

Abstract
Nonlinear filtering problems are encountered in many applications, and one solution approach is the extended Kalman filter, which is not always convergent. Therefore, it is crucial to identify conditions under which the extended Kalman filter provides accurate approximations. This paper generalizes two significant results from picard1991efficiency on the efficiency of the continuous-time extended Kalman filter to a more general setting where the observation noise may be state-dependent but does not allow signal reconstruction from the quadratic variation of the observation process as in epidemic models. Firstly, we show that when the observation’s drift coefficient is strongly injective and the signal’s and observation’s drift become nearly linear for the diffusion scaling coefficient $\epsilon\to 0$, the estimation error is of order $\sqrt{\epsilon}$. Subsequently, we establish conditions under which the impact of the initial filtering error decays exponentially fast.
1 Introduction
This research investigates a nonlinear filtering problem in the presence of state-dependent noise. Many physical phenomena are best described by nonlinear sto-
chastic dynamics, in which the diffusion coefficient depends on the system’s state. These characteristics are essential for accurately modeling and understanding such systems. Therefore, filtering problems associated with these phenomena are nonstandard and necessitate alternative methods to the standard Kalman filtering approach. For instance, estimating the angular procession of a rotating spacecraft requires precise nonlinear filtering techniques to account for the state-dependent variations in noise. Similarly, the design of phase-locked loops, which are essential in communication systems to maintain signal synchronization, also requires these advanced filtering methods 1099828. These examples highlight the practical significance and wide-ranging applications of our study. The model that we will be interested in is the following.
$\begin{split}dY(t)&=f(t,Y(t),\pazocal{Z}(t))dt+\sqrt{\varepsilon}\sigma(t,Y(t),\pazocal{Z}(t))dW^{(1)}(t)+\sqrt{\varepsilon}g(t,Y(t),\pazocal{Z}(t))dW^{(2)}(t),\\[2.15277pt] dZ(t)&=h(t,Y(t),\pazocal{Z}(t))dt+\sqrt{\varepsilon}\ell(t,Y(t),\pazocal{Z}(t))dW^{(2)}(t),\end{split}$ (1)
with $(Y(0),Z(0))=(y_{0},z_{0})$, and where $Y(t)\in\mathbb{R}^{n}$ is our hidden signal and $Z(t)\in\mathbb{R}^{d}$ is our observation process at time $t\geq 0$. Additionally, $\pazocal{Z}(t)=\{Z_{s},s\leq t\}$ denotes the path of the observation process until the time point $t$. Finally, $\varepsilon$ is a small positive constant and $W^{(i)},~i=1,2$ are two independent standard Wiener processes in $\mathbb{R}^{d_{1}}$and $\mathbb{R}^{d_{2}}$ respectively (with $n,d,d_{1},d_{2}$ being positive integers).
This model is quite similar to the model in picard1991efficiency. However, we address a more general setting by allowing the diffusion coefficient of the observation process to be non-constant and state-dependent. For such a model, when the quadratic variation of the observation process is informative and can be computed in practice, it can be possible to reconstruct the signal under some injectivity conditions. This reconstruction is often not possible as in the context of epidemic models for which injectivity conditions are not fulfilled. It is therefore important to know whether the extended Kalman filter is efficient for such filtering problems.
Literature review
Many papers in the literature have been devoted to this type of filtering problem for which the diffusion coefficient of the observation is state-dependent. For instance, mclane1969optimal have proposed an optimal linear filter by restricting the filter structure to the linear class with respect to the observation. takeuchi1981least; takeuchi1985least have provided a more accurate approximation for the least-square optimal estimate as well as the convergence property of that estimate. More recently, joannides1995nonlinear; joannides1997nonlinear have considered nonlinear filtering with perfect observation and noninformative quadratic variation, which helps design robust and efficient algorithms in the case of filtering with small observation noise. Moreover, the case of state-dependent and small noise for Markovian diffusion has been investigated by picard1993estimation and he suggested a stochastic algorithm for the approximation of the optimal estimate.
On the other hand, several authors have considered the nonlinear filtering problem with small noise. Most of them are interested in finding efficient asymptotic approximate filters katzur1984asymptotic1; katzur1984asymptotic2. picard1991efficiency proved some convergence results on the extended Kalman filter for filtering problems with constant and small diffusion coefficient in the observation process dynamic.
Our contribution
The main contributions of this paper are the generalization of two important results on error estimates of the extended Kalman filter from picard1991efficiency. First, we prove that the normalized the filtering error is of order $\sqrt{\varepsilon}$ for the more general model 1 under some assumptions. Second, we show that the filtering error is bounded over time, and the impact of the initial error on the filtering error vanishes exponentially fast.
Paper outline
In Section 2, we introduce a motivating example from mathematical epidemiology, then Section 3 describes the derivation of the extended Kalman filter for (1), and presents some relevant definitions related to dynamical systems. Section 4 focuses on the estimation error of the extended Kalman filter, and we prove that under some assumptions, the filtering error is of order $\sqrt{\varepsilon}$. Finally, in Section 5, we show that the filtering error is bounded, and the initial error is forgotten exponentially fast when certain detectability conditions are met. The Appendix collects auxiliary results needed for the proofs of our theorems.
2 Motivating example: $SI^{\pm}S$ epidemic model
In this section, we introduce a filtering system from mathematical epidemiology that is related to stochastic epidemic models with partial information and can be reformulated using appropriate time scaling as the filter system specified in (1). Let us consider the $SI^{\pm}S$ epidemic model in which a population is divided into three compartments according to the health state of individuals: susceptible ($S$), undetected infected ($I^{-}$), and detected infected ($I^{+}$). The associated flowchart is depicted in Figure 1.
Figure 1: Flowchart of the $SI^{\pm}$ compartmental model.
This model assumes that when a susceptible individual is infected, he or she is first presymptomatic or asymptomatic. This individual is accounted for in the $I^{-}$ compartment. After a test that occurs at a rate of $\alpha$, the undetected infected individuals become detected and move to the $I^{+}$ compartment. Here, people are in quarantine and can no longer infect new susceptible individuals. Both detected and undetected infected individuals recover from the infection and return to the susceptible compartment. The parameters $\beta$, $\rho^{-}$, and $\rho^{+}$ are the infection intensity, the recovery rates for the undetected and detected infected, respectively. We consider the population size $N$ to be constant, that is, $S+I^{-}+I^{+}=N$. A more detailed description of such models can be found in kamkumo2025estimating; kamkumo2025stochastic; ouabo2025phd, njiasse2025stoch; mbouandi2025phd. The stochastic dynamics of such a compartmental model can be derived using a continuous-time Markov chain approach as described in BrittonPardoux2019 and in kamkumo2025stochastic. For large population sizes $N$, diffusion approximations can be derived for the state vector $X$, which contains the relative subpopulation sizes. These are stochastic differential equations of the form
$dX(t)$ $=\overline{F}(X(t))dt~+~\dfrac{1}{\sqrt{N}}\overline{\sigma}(X(t))dB(t),~~X(0)=x_{0},$ (2)
where coefficients $\overline{F}$, $\overline{\sigma}$ are non-linear functions and $B$ is a vector of independent standard Brownian motions. For instance, in the case of the above $SI^{\pm}S$ epidemic model, if we work with the reduced state $X=(X_{1},X_{2})^{\top}=(\overline{I}^{-},\overline{I}^{+})^{\top}=N^{-1}({I}^{-},{I}^{+})^{\top}$, Since we assume a constant population size, the proportion of susceptible individuals can be removed from the state, as it can be obtained from $\overline{S}=1-\overline{I}^{+}-\overline{I}^{-}$). The expressions of the functions $\overline{F}$ and $\overline{\sigma}$ read as follows, when the state $X$ is split into hidden component $Y=\overline{I}^{-}$ and observable component $Z=\overline{I}^{+}$:
$\overline{F}\left(\begin{array}[]{l}y\\ z\end{array}\right)$ $=\left(\begin{array}[]{l}\beta(1-y-z)y-(\alpha+\rho^{-})y\\[4.30554pt] \alpha y-\rho^{+}z\end{array}\right),$ (7)
[…]
and $B=\left({B^{(1)}},\ldots,{B^{(4)}}\right)^{\top}$.
Let $\varepsilon={1}/{\sqrt{N}}$, $\tau={t}/{\varepsilon}=\sqrt{N}t$ and $U^{\varepsilon}:=\varepsilon^{-1}\left(X(\varepsilon\tau)-x_{0}\right)$. Using Equation (2) to derive the dynamics of the state $U^{\varepsilon}$ under the time scaling $t=\varepsilon\tau$, we get
$dU^{\varepsilon}(\tau)$ $=\overline{F}(x_{0}+\varepsilon U^{\varepsilon}(\tau))d\tau~+~\sqrt{\varepsilon}\overline{\sigma}(x_{0}+\varepsilon U^{\varepsilon}(\tau))dW(\tau),~~U^{\varepsilon}(0)=0,$ (14)
where $W(\tau)=\varepsilon^{-1/2}B(\varepsilon\tau)$. The scaling property of Brownian motion implies that the components of $W$ are independent standard Brownian motions. On a small time window with length depending on $\varepsilon$, functions $\overline{F}$ and $\overline{\sigma}$ can be approximated by linear functions since in the timescale $\tau$ their respective nonlinearities are less pronounced. The filtering system obtained for the components of $U^{\varepsilon}=(Y,Z)^{\top}$ with hidden signal $Y$ and observation $Z$ has exactly the structure of the system (1), with following functions coefficients:
$f(t,y,z)~=$ $~\beta(1-y-z)y-(\alpha+\rho^{-})y,\quad h(t,y,z)~=~\alpha y-\rho^{+}z,$ (15)
$\sigma(t,y,z)~=$ $~\left(\begin{array}[]{cc}\sqrt{\beta(1-y-z)y},&-\sqrt{\rho^{-}y}\end{array}\right),\quad g(t,y,z)~=~\left(\begin{array}[]{cc}-\sqrt{\alpha y},&0\end{array}\right),$ (18)
$\ell(t,y,z)~=$ $~\left(\begin{array}[]{cc}\sqrt{\alpha y},&-\sqrt{\rho^{+}z}\end{array}\right),$ (20)
with $W^{(1)}=(W_{1},W_{2})^{\top}$ and $W^{(2)}=(W_{3},W_{4})^{\top}$.
3 Extended Kalman filter and some general concepts
3.1 Extended Kalman filter
The Kalman filtering procedure is well known and has been successfully applied in many domains. However, it requires the filtering problem to be linear. Some attempts have been made to apply the Kalman filter core idea to nonlinear models. One of the most prominent attempts is the so-called extended Kalman filter. This extended approach was first introduced by gelb1974applied, and an empirical justification was proposed by pardoux1991filtrage. This approach is described in many books, for instance bain2008fundamentals for more details. In the following computations and the rest of this work, functions will be of three variables: time $t$, signal $Y(t)$ and observation path history $\pazocal{Z}(t)$. However, for the sake of short notation, we will in some expressions omit dependence on the observation path history. Let us revisit the main step of the derivation of the extended Kalman filter for the nonlinear filtering problem (1). Let us assume that $\overline{Y}$ is the solution of the associated ordinary differential equation to our signal stochastic differential equation obtained by removing the diffusion term,
$\dfrac{d\overline{Y}(t)}{dt}=f(t,\overline{Y}(t)),~~~~\overline{Y}(0)=y_{0}.$ (21)
On a small interval of time $0,\Delta[$, with $\Delta>0$, we can perform the Taylor-like expansion of our nonlinear filtering system drift around the solution $\overline{Y}$, and replace $Y$ by $\overline{Y}$ in the diffusion coefficients. Hence, we obtain the linearized filtering system
[…]
where $\nabla_{y}f$ and $\nabla_{y}h$ are the derivatives of $f$ and $h$ with respect to the signal argument, $Y$ respectively. Since the latter system is linear in $Y$, one can argue that in a small interval of time the equations satisfied by the original signal $Y$ and observation $Z$ are nearly linear. This setting is then suitable for the result in LiptserShiryaevVolII2001 on conditionally Gaussian process. By applying it, we compute from the linearized problem the approximations of the conditional mean $\widehat{Y}$ and the conditional variance $Q^{\varepsilon}$ that satisfy
$\widehat{Y}(t)~=$ $\widehat{Y}(0)+\int_{0}^{t}\left(\nabla_{y}f(s,\overline{Y}(s))(\widehat{Y}(s)-\overline{Y}(s))+f(s,\overline{Y}(s))\right)ds$
$+\int_{0}^{t}G(s)\big[dZ(s)-(h(s,\widehat{Y}(s))-\nabla_{y}h(s,\overline{Y}(s))\overline{Y}(s)+\nabla_{y}h(s,\overline{Y}(s))\widehat{Y}(s))ds\big],$ (22)
where
$G(t)=\big[\varepsilon g\ell^{\top}(t,\overline{Y}(t))+Q^{\varepsilon}(t)\nabla_{y}h^{\top}(t,\overline{Y}(t))\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}(t,\overline{Y}(t)),$ (23)
and
$\dot{Q^{\varepsilon}}(t)~=$ $-\big[\varepsilon g\ell^{\top}+Q^{\varepsilon}(t)\nabla_{y}h^{\top}\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}\big[\varepsilon g\ell^{\top}+Q^{\varepsilon}(t)\nabla_{y}h^{\top}\big]^{\top}(t,\overline{Y}(t))$
$+\nabla_{y}f(t,\overline{Y}(t))Q^{\varepsilon}(t)+Q^{\varepsilon}(t)\nabla_{y}f^{\top}(t,\overline{Y}(t))+\varepsilon(\sigma\sigma^{\top}+gg^{\top})(t,\overline{Y}(t)).$ (24)
By this method, it is possible to construct a mapping which relates any observable process of linearization such as $\overline{Y}$ to the corresponding conditional mean of the filter. Then, the extended Kalman filter by definition is the fixed point for this mapping, see bain2008fundamentals. Hence, we deduce that the extended Kalman filter for the nonlinear filtering problem (1) is given by the conditional mean $M$ that satisfies
$M(t)~=$ $M(0)+\int_{0}^{t}f(s,M(s))ds+\int_{0}^{t}G(s)\big[dZ(s)-h(s,M(s))ds\big],$ (25)
where
$G(t)=\big[\varepsilon g\ell^{\top}(t,M(t))+Q^{\varepsilon}(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}(t,M(t)),$ (26)
and the conditional covariance matrix with dynamics
$\dot{Q^{\varepsilon}}(t)=$ $~-\big[\varepsilon g\ell^{\top}+Q^{\varepsilon}(t)\nabla_{y}h^{\top}\big]\big(\varepsilon\ell\ell^{\top}\big)^{-1}\big[\varepsilon g\ell^{\top}+Q^{\varepsilon}(t)\nabla_{y}h^{\top}\big]^{\top}(t,M(t))$ (27)
$~+\nabla_{y}f(t,M(t))Q^{\varepsilon}(t)+Q^{\varepsilon}(t)\nabla_{y}f^{\top}(t,M(t))+\varepsilon(\sigma\sigma^{\top}+gg^{\top})(t,M(t)).$ (28)
We can now define what will be called in the sequel, the extended Kalman filter for the filtering system (1). For the sake of simpler equations, we introduce the scaled covariance matrix $Q=\dfrac{1}{\varepsilon}Q^{\varepsilon}$.
Definition 3.1
Let $(M(t))_{t\geq 0}$ and $(Q^{\varepsilon}(t))_{t\geq 0}$ be two observable processes with values respectively in $\mathbb{R}^{n}$ and the set of positive-definite symmetric matrices of order $n$. A process which at any time $t$ is Gaussian with mean $M(t)$ and covariance matrix $Q^{\varepsilon}(t)=\varepsilon Q(t)$ will be said to be an extended Kalman filter for (1) if $M(t)$ is solution of
$M(t)=M(0)+\int_{0}^{t}f(s,M(s))ds+\int_{0}^{t}G(s)\big[dZ(s)-h(s,M(s))ds\big],$ (29)
where
$G(t)=\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}(t,M(t)),$ (30)
and $Q(t)$ is solution of the Riccati equation
$\dot{Q}(t)$ $=\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big](t,M(t))Q(t)+Q(t)\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big]^{\top}(t,M(t))$
$~~~~~~-Q(t)[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h](t,M(t))Q(t)+a(t,M(t)),$ (31)
where
$a(t,M(t))=\sigma\sigma^{\top}(t,M(t))+g(I-\ell^{\top}(\ell\ell^{\top})^{-1}\ell)g^{\top}(t,M(t)).$ (32)
3.2 Some definitions
We will now introduce some essential concepts for this study. We will use the expression a "family of" function or process when this function or process depends on $\varepsilon$. Furthermore, we fix a family of filtered probability spaces, denoted by $\left(\Omega,\pazocal{F},\mathbb{F},\mathbb{P}\right)$ and we assume that the family of functions $f,~h,~\sigma,~g,$$\ell$ are measurable and locally bounded. The filtration, denoted by $\mathbb{F}$, is such that all the processes defined in the paper are adapted with respect to it.
Definition 3.2(Observable Process)
Let us denote by $\mathbb{F}^{Z}$ the observable filtration generated by the observation process $(Z(t))_{t\geq 0}$. We will call an observable process any $\mathbb{F}^{Z}$-predictable process.
This means that any observable process is adapted with respect to the natural filtration $\mathbb{F}^{Z}$.
Definition 3.3
Let $\varphi:[0,\infty)\times\mathbb{R}^{n}\to\mathbb{R}^{n},$ be a family of functions and $x_{1},x_{2}\in\mathbb{R}^{n}$.
  1. 1.
    $\varphi$ is said to be almost linear if there exists a family of matrix-valued observable processes $F:[0,\infty)\to\mathbb{R}^{n\times n}$ such that
    $|\varphi(t,x_{1})-\varphi(t,x_{2})-F(t)(x_{1}-x_{2})|\leq\mu^{\varepsilon}|x_{1}-x_{2}|,$ (33)
    for some family of numbers $\mu^{\varepsilon}$ converging to zero as $\varepsilon\to 0$; the process $F$ will be called an almost derivative of $f$.
  2. 2.
    The function $\varphi$ will be said to be strongly injective if
    $|\varphi(t,x_{1})-\varphi(t,x_{2})|\geq c|x_{1}-x_{2}|,$ (34)
    for some $c>0$.
Definition 3.4
Let $L^{q}$ be the space of random variables with finite moment of order $q\geq 1$. The norm of a random variable $X$ in $L^{q},~q\geq 1$, is denoted by $\left|X\right|_{q}$ and is defined by $\left|X\right|_{q}=\Big(\mathbb{E}[|X|^{q}]\Big)^{1/q}$.
Definition 3.5
A family of process $(\pazocal{X}^{\varepsilon}(t))_{t\geq 0}$ is said to be bounded in $L^{\infty-}$ if for any $q\in(0,\infty)$, there exists $\varepsilon_{q}>0$ such that $|\pazocal{X}^{\varepsilon}(t)|_{q}$ is bounded uniformly in $t\geq 0$ and $0<\varepsilon<\varepsilon_{q}$.
Definition 3.6(Stochastic Process of Order $\varepsilon^{\alpha}$)
A family of stochastic processes $(\pazocal{X}^{\varepsilon}(t))_{t\geq 0}$, $\varepsilon>0$ is said to be of order $\varepsilon^{\alpha}$ for some $\alpha>0$ if for any $q\in[1,\infty)$ there exist constants $\varepsilon_{q}>0,~C_{q}>0$, such that
$\varepsilon^{-\alpha}\left|\pazocal{X}^{\varepsilon}(t)\right|_{q}\leq C_{q}~~\text{for all}~~(t,\varepsilon)\in[0,\infty)\times(0,\varepsilon_{q}).$
Similarly, a family of random variables $\pazocal{X}^{\varepsilon}$, $\varepsilon>0$ is said to be of order $\varepsilon^{\alpha}$ for some $\alpha>0$ if for any $q\in[1,\infty)$ there exist constants $\varepsilon_{q}>0,~C_{q}>0$, such that
$\varepsilon^{-\alpha}\left|\pazocal{X}^{\varepsilon}\right|_{q}\leq C_{q}~~\text{for all}~~\varepsilon\in(0,\varepsilon_{q}).$
In the study of dynamical systems, stability properties are powerful to characterize the qualitative behavior of a solution. We introduce two concepts of stability for a family of matrices.
Definition 3.7
Let $(A(t))_{t\geq 0}$ be a family of measurable locally bounded processes with values in the class of $n\times n$ matrices, let $\zeta$ be the matrix-valued solution of
$\dot{\zeta}(t)=A(t)\zeta(t),~~~\zeta(0)=I.$ (35)
We will say that $(A(t))_{t\geq 0}$ is exponentially stable if there exist some positive constant numbers $C$ and $c$ such that for $s\leq t$,
$|\zeta(t)|\leq Ce^{(-c(t-s))}|\zeta(s)|.$ (36)
Remark 3.8
In case of constant matrix $A$, we can show that $A$ is exponentially stable if and only if its eigenvalues have negative real part.
Definition 3.9
Let $A,B\in\pazocal{M}_{n}$, where $\pazocal{M}_{n}$ denotes the class of $n\times n$ real matrices.
We write $A\succeq B$ if $A$ and $B$ are symmetric and ABA-B is positive semi-definite.
We write $A\succ B$ if $A$ and $B$ are symmetric and ABA-B is positive definite.
Definition 3.10
Consider a family $(\pazocal{K}(t))_{t\geq 0}$ of absolutely continuous adapted (w.r.t. the observable filtration) processes with values in the class of symmetric positive definite $n\times n$ matrices and a family $(k(t))_{t\geq 0}$ of locally bounded deterministic functions with positive values. We will say that a family $(A(t))_{t\geq 0}$ of measurable locally bounded process with values in the class $n\times n$ matrices, is $(\pazocal{K}(t),k(t))_{t\geq 0}$-stable if for all $t\geq 0$,
$\dot{\pazocal{K}}(t)\succeq A(t)\pazocal{K}(t)+\pazocal{K}(t)A^{\top}(t)+k(t)\pazocal{K}(t).$ (37)
Remark 3.11
The link between these two stability concepts is given below by
Lemma 5.3.
4 Error estimation for the extended Kalman filter
When estimating a hidden signal, it is of relevance to know whether the scale of the error between the estimate and the true signal is large or not. This is why we are interested in conditions under which the extended Kalman filter for problem (1) provides a good approximation in the case of a strongly injective drift coefficient of the observation process. We show in the following theorem that the strongly injectivity combined with an initial guess of order $\sqrt{\varepsilon}$ and a set of technical conditions ensure that the estimation error Y(t)M(t)Y(t)-M(t) is of order $\sqrt{\varepsilon}$ at any future time point $t$.
Theorem 4.1
Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be an extended Kalman filter for problem (1) such that $M$ and $Q=\frac{1}{\varepsilon}Q^{\varepsilon}$ satisfy (29) and (31) respectively, and in addition
  1. (A1)
    The variable $Q^{-1/2}(0)(Y(0)-M(0))$ is of order $\sqrt{\varepsilon}$.
  2. (A2)
    The functions $\sigma$, $g$ and $\ell$ are bounded.
  3. (A3)
    The function $f$ and $h$ are $\pazocal{C}^{1}$ and almost linear with respect to the signal.
  4. (A4)
    The function $h$ is strongly injective with respect to the signal,
    $a=\sigma\sigma^{\top}-g(I-\ell^{\top}(\ell\ell^{\top})^{-1}\ell)g^{\top}$ and $(\ell\ell^{\top})^{-1}$ are uniformly elliptic,
    and the quotient between the largest and the smallest eigenvalues of $Q(0)$ is bounded.
Then $Q^{-1/2}(t)(Y(t)-M(t))$ is of order $\sqrt{\varepsilon}$.
Remark 4.2
Assumptions (A2)-(A4) of this theorem can be checked for the $SI^{\pm}S$ epidemic model with constant population size. Moreover, all compartment sizes are assumed strictly positive. The assumption (A2) is fulfilled because of the strictly positive compartment sizes. The hidden signal and observation drift coefficients are respectively,
$f(Y(\tau))=$ $~~\beta(1-(X_{1}(0)+\varepsilon Y(\tau))-(X_{2}(0)+\varepsilon Z(\tau)))(X_{1}(0)+\varepsilon Y(\tau))$-
$~~(\alpha+\rho^{-})(X_{1}(0)+\varepsilon Y(\tau)),$
$h(Y(\tau))=$ $~~\alpha(X_{1}(0)+\varepsilon Y(\tau))-\rho^{+}(X_{1}(0)+\varepsilon Y(\tau)).$
These functions are $\pazocal{C}^{1}$ in $Y$ and $h$ is almost linear since it is linear, hence (A3) is fulfilled. The function $f$ is also almost linear with almost derivative the coefficient of the linear term in $f$, which is
$\beta\varepsilon\left(1-2X_{1}(0)-(X_{2}(0)+\varepsilon Z(\tau))\right)-(\alpha+\rho^{-})\varepsilon.$
For assumption (A4), $h$ is strongly injective since it is linear, and $a$ and $(\ell\ell^{\top})^{-1}$ are strictly positive scalars since all compartments sizes are strictly positive. The initial matrix $Q(0)$ is a scalar and can be chosen to satisfy the assumption.
For the first assumption, the scaled initial error estimate $Q^{-1/2}(0)(Y(0)-M(0))$ should be of order $\sqrt{\varepsilon}$. This implies that the accuracy of the scaled initial estimate of absolute population size should be of order $\sqrt[4]{N}$ since $\varepsilon=\dfrac{1}{\sqrt{N}}$.
In order to prove Theorem 4.1, we first establish the following lemma and then we will show that the theorem is a particular case of this lemma.
Lemma 4.3
Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be an EKF satisfying assumptions (A1)-(A3) of Theorem 4.1. Suppose also that
$Q(t)\big(\nabla_{y}h^{\top}\nabla_{y}h\big)(t,M(t))Q(t)+a(t,M(t))\succeq k(t)Q(t),$ (38)
for some family of deterministic positive functions $k(t)$ such that
$Q(t)+Q^{-1}(t)\preceq Ck(t)I.$ (39)
Then $Q^{-\frac{1}{2}}(t)\big(Y(t)-M(t)\big)$ is of order $\sqrt{\varepsilon}$.
The proof of this Lemma is based on the following lemma, which is proven in picard1991efficiency.
Lemma 4.4
Let $\widetilde{W}(t)$ be a $\mathbb{F}$-Brownian motion with values in $\mathbb{R}^{r}$, $n\in\mathbb{N}$, and let $A(t)$ be a family of $\mathbb{F}$- adapted $(\pazocal{K}(t),k(t))$-stable processes where $\pazocal{K}(t)$ is $\mathbb{F}$ adapted. We suppose that $\pazocal{X}(t)$ is a family of $\mathbb{R}^{n}$-valued semi-martingales satisfying
$d\pazocal{X}(t)=A(t)\pazocal{X}(t)dt+F(t)dt+D(t)d\widetilde{W}(t),$ (40)
where $F(t)$ and $D(t)$ are predictable processes satisfying
$\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)F(t)$ $\leq~\alpha k(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)+\pazocal{O}(k(t)),$ (41)
$D^{\top}(t)\pazocal{K}^{-1}(t)D(t)$ $=~\pazocal{O}(k(t)),$ (42)
for some $\alpha\in(0,\frac{1}{2})$.
If $\pazocal{X}^{\top}(0)\pazocal{K}^{-1}(0)\pazocal{X}(0)$ is bounded in $L^{q}$, then for any $q\geq 1$ and $\varepsilon$ small enough, the process $\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)$ is bounded in $L^{q}$.
Proof
of Lemma 4.3
Let us consider $\pazocal{K}(t)=\varepsilon Q(t)$ and let us recall the dynamics of our signal $Y$ and of the extended Kalman filter $M$ from (1) and (29), respectively.
$dY(t)~=$ $f(t,Y(t))dt+\sqrt{\varepsilon}\sigma(t,Y(t))dW^{(1)}(t)+\sqrt{\varepsilon}g(t,Y(t))dW^{(2)}(t),$
$dM(t)~=$ $f(t,M(t))dt+\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\times\big[dZ(t)-h(t,M(t))dt\big],$
$~=$ $f(t,M(t))dt+\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\times$
$\big[h(t,Y(t))dt+\sqrt{\varepsilon}\ell(t,Y(t))dW^{(2)}(t)-h(t,M(t))dt\big].$
Thus, if we denote the difference YMY-M by $\pazocal{X}$, its dynamics is given by
$d\pazocal{X}(t)~=$ $\big[f(t,Y(t))-f(t,M(t))\big]dt$
$-\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\times\big[h(t,Y(t))-h(t,M(t))\big]dt$
$+\sqrt{\varepsilon}\sigma(t,Y(t))dW^{(1)}(t)+\sqrt{\varepsilon}g(t,Y(t))dW^{(2)}(t)$
$-\sqrt{\varepsilon}\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\ell(t,Y(t))dW^{(2)}(t)$
$~=$ $\big[f(t,Y(t))-f(t,M(t))\big]dt$
$-\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\times\big[h(t,Y(t))-h(t,M(t))\big]dt$
$+\sqrt{\varepsilon}\sigma(t,Y(t))dW^{(1)}(t)+\sqrt{\varepsilon}\Big[g(t,Y(t))$
$-\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\ell(t,Y(t))\Big]dW^{(2)}(t).$
Here, the second equality is obtained by simple expansion. Adding and subtracting the same term and rearranging, we get
$d\pazocal{X}(t)~=$ $\Big[f(t,Y(t))-f(t,M(t))-\Big(\nabla_{y}f(t,M(t))-\nabla_{y}f(t,M(t))\Big)\pazocal{X}(t)\Big]dt$
$-\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}$
$\times\big[h(t,Y(t))-h(t,M(t))-\big(\nabla_{y}h(t,M(t))-\nabla_{y}h(t,M(t))\big)\pazocal{X}(t)\big]dt$
$+\sqrt{\varepsilon}\Big(\sigma(t,Y(t)),g(t,Y(t))$-
$\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\ell(t,Y(t))\Big)\left(\begin{array}[]{c}dW^{(1)}(t)\\ dW^{(2)}(t)\end{array}\right)$
$~=$ $\left[\nabla_{y}f(t,M(t))-\big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}\big)(\ell\ell^{\top})^{-1}\nabla_{y}h(t,M(t))\right]\pazocal{X}(t)dt$
$+\left[f(t,Y(t))-f(t,M(t))-\nabla_{y}f(t,M(t))\pazocal{X}(t)\right$.
$\left.-\Big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\Big)(\ell\ell^{\top})^{-1}\right$.
$\times\left.\Big(h(t,Y(t))-(h(t,M(t))-\nabla_{y}h(t,M(t))\pazocal{X}(t))\Big)\right]dt+\sqrt{\varepsilon}\Big(\sigma(t,Y(t)),$
$g(t,Y(t))-\big[g\ell^{\top}+Q(t)\nabla_{y}h^{\top}\big](t,M(t))\big(\ell\ell^{\top}\big)^{-1}\ell(t,Y(t))\Big)\left(\begin{array}[]{c}dW^{(1)}(t)\\ dW^{(2)}(t)\end{array}\right).$
We are then in the setting of Lemma 4.4 with
$d\pazocal{X}(t)~=$ $A(t)\pazocal{X}(t)dt+F(t)dt+D(t)d\widetilde{W}(t),$
where
$A(t)~:=$ $\nabla_{y}f(t,M(t))-\big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}\big)(\ell\ell^{\top})^{-1}\nabla_{y}h(t,M(t)),$
$F(t)~:=$ $\left[f(t,Y(t))-f(t,M(t))-\nabla_{y}f(t,M(t))\pazocal{X}(t)-\Big(g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\Big)\right$.
$\times(\ell\ell^{\top})^{-1}(t,M(t))\left.\Big(h(t,Y(t))-(h(t,M(t))-\nabla_{y}h(t,M(t))\pazocal{X}(t))\Big)\right],$
$D(t)~:=$ $\sqrt{\varepsilon}\Big(\sigma(t,Y(t)),~g(t,Y(t))-\big[g\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\big]\big(\ell\ell^{\top}\big)^{-1}\ell(t,Y(t))\Big),$
$\widetilde{W}(t):=$~ $\left(\begin{array}[]{c}W^{(1)}(t)\\ W^{(2)}(t)\end{array}\right).$
By definition, for any $k(t)$, saying that $A(t)$ is $(Q(t),k(t))$-stable means that
$\dot{Q}(t)$ $=\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]Q(t)+Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]$
$~~~~~-Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+a(t,M(t))$
$\succeq\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]Q(t)-Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)$
$~~~~~+Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}-Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+k(t)Q(t),$
and this implies that $a(t,M(t))\succeq-Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+k(t)Q(t)$, which is equivalent to the condition (38). Now, we have to check the two conditions (41) and (42).
Condition (42) can be deduced from the assumption (39) as follows. We have that
$Q(t)+Q^{-1}(t)\preceq Ck(t)I$ This implies that
$D^{\top}(t)(Q(t)+Q^{-1}(t))D(t)$ $\preceq D^{\top}(t)Ck(t)ID(t)$ (43)
$\text{and thus }~~D^{\top}(t)Q(t)D(t)+D^{\top}(t)Q^{-1}(t)D(t)$ $\preceq Ck(t)|D(t)|^{2}.$ (44)
Since $Q(t)$ and $Q^{-1}(t)$ are positive definite, hence $D^{\top}(t)Q(t)D(t)\succeq 0$ and
$D^{\top}(t)Q^{-1}(t)D(t)\succeq 0$. In addition, $|D(t)|$ is bounded, then $D^{\top}(t)Q^{-1}(t)D(t)=O(k(t))$.
Now, let us prove the first condition. Instead of the first condition, we are going to show that
$F^{\top}(t)\pazocal{K}^{-1}(t)F(t)\leq\alpha_{\epsilon}k^{2}(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t),$ (45)
which will imply that
$\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)F(t)\leq\alpha^{1/2}_{\epsilon}k(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t),$ (46)
according to the Remark 4.5 below.
We have
$F^{\top}(t)Q^{-1}(t)F(t)$ $=\Big[B-\left(g(t,M(t))\ell^{\top}(t,M(t))-Q(t)\nabla_{y}h^{\top}\right)(\ell\ell^{\top})^{-1}H\Big]^{\top}$
$~~~~~\times Q^{-1}(t)\times\Big[B-\left(g(t,M(t))\ell^{\top}(t,M(t))-Q(t)\nabla_{y}h^{\top}\right)(\ell\ell^{\top})^{-1}H\Big],$
with
$B$ $:=f(t,Y(t))-f(t,M(t))-\nabla_{y}f(t,M(t))(Y(t)-M(t)),$
$H$ $:=h(t,Y(t))-h(t,M(t))-\nabla_{y}h(t,M(t))(Y(t)-M(t)).$
Thus, since $f$ and $h$ are almost linear, we obtain the following estimate,
$F^{\top}(t)Q^{-1}(t)F(t)=$ $~B^{\top}Q^{-1}B-H^{\top}(\ell\ell^{\top})^{-1}(g(t,M(t))\ell^{\top}$
$+Q(t)\nabla_{y}h(t,M(t)))^{\top}(\ell\ell^{\top})^{-1}Q^{-1}B$
$+H^{\top}(\ell\ell^{\top})^{-1}\Big(g(t,M(t))\ell^{\top}+Q(t)\nabla_{y}h(t,M(t))\Big)^{\top}(\ell\ell^{\top})^{-1}Q^{-1}$
$\times\Big(g(t,M(t))\ell^{\top}+Q(t)\nabla_{y}h(t,M(t))\Big)(\ell\ell^{\top})^{-1}H(t)$
$-B^{\top}Q^{-1}\Big(g(t,M(t))\ell^{\top}+Q(t)\nabla_{y}h(t,M(t))\Big)^{\top}(\ell\ell^{\top})^{-1}H(t)$
$\leq$ $~\mu_{\varepsilon}(1+|Q(t)|+|Q^{-1}(t)|)|Y(t)-M(t)|^{2},$
because $g$, $\ell$ and $(\ell\ell^{\top})^{-1}$ are bounded.
In addition, $Q(t)+Q^{-1}(t)\preceq Ck(t)I$ implies that $|Q(t)|+|Q^{-1}(t)|\leq C|k(t)|$, therefore we can rewrite the last inequality as
$F^{\top}(t)Q^{-1}(t)F(t)\leq\mu_{\varepsilon}Ck(t)(Y(t)-M(t))^{\top}I(Y(t)-M(t)).$ (47)
And then, from $Q(t)\preceq ck(t)I$, we obtain $Ck(t)Q^{-1}\succeq I$ and the inequality (47) can take the form
$F^{\top}(t)Q^{-1}(t)F(t)\leq\mu_{\varepsilon}Ck^{2}(t)(Y(t)-M(t))^{\top}Q^{-1}(t)(Y(t)-M(t)).$ (48)
The condition (45) is satisfied for some $\mu_{\varepsilon}\longrightarrow 0$. Therefore, the conclusion of lemma 4.4 applies, and we deduce that $Q^{-\frac{1}{2}}(t)(Y(t)-M(t))$ is of order $\sqrt{\varepsilon}$. ∎
Remark 4.5
This remark allows us to use condition (45) instead of condition (41). We use the Cauchy-Schwarz inequality corresponding to the inner product defined by the positive definite matrix $\pazocal{K}^{-1}(t)$.
$\left|\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)F(t)\right|$ $\leq\Big(\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)\Big)^{\frac{1}{2}}\Big(F^{\top}(t)\pazocal{K}^{-1}(t)F(t)\Big)^{\frac{1}{2}}$
$\leq\Big(\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)\Big)^{\frac{1}{2}}\Big(\alpha_{\epsilon}k^{2}(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)\Big)^{\frac{1}{2}}$
$\leq\alpha^{1/2}_{\epsilon}k(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t).$
Remark 4.6
Note that in the proof above, for the sake of clarity we use $A(t)$ is $(Q(t),k(t))$-stable instead of $A(t)$ is $(\varepsilon Q(t),k(t))$-stable since both are equivalent.
Proof
of Theorem 4.1.
Since we have assumed that $a$ and $(\ell\ell^{\top})^{-1}$ are uniformly elliptic and that $h$ is strongly injective, we have
$Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+a(t,M(t))\succeq c(Q^{2}(t)+I)\succeq c\left(Q(t)+Q^{-1}(t)\right)Q(t).$ (49)
If there exists a family of deterministic positive functions $p(t)$ such that $p^{-1}(t)Q(t)$ is uniformly bounded and elliptic, then the conditions of Lemma 4.3 will be satisfied with $k(t)$ proportional to $p(t)+p^{-1}(t)$. Namely, if we assume that $p^{-1}(t)Q(t)$ is uniformly bounded and elliptic, it implies that there exist two positive real constants $b_{1}$ and $b_{2}$ satisfying
$b_{1}I\preceq p^{-1}(t)Q(t)\preceq b_{2}I.$ (50)
From the first part of this inequality we deduce that $b_{1}p(t)I\preceq Q(t)$ and from the second part we get $\dfrac{p^{-1}(t)}{b_{2}}I\preceq Q(t)^{-1}$. Combining these two inequalities, we get
$Q(t)+Q(t)^{-1}\succeq\left(b_{1}p(t)+\dfrac{p^{-1}(t)}{b_{2}}\right)I\succeq b\left(p(t)+p^{-1}(t)\right)I,$ (51)
for some constant $b$. Thus, we only have to prove the existence of such a family $p(t)$.
Let $p(0)$ be the trace of $Q(0)$, as a reminder the dynamics of $Q$ is given by
$\hskip-2.84544pt\dot{Q}(t)=$ $~\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))Q(t)+Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))$ (52)
$~-Q(t)\Big[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))Q(t)+a(t,M(t)).$
Then, applying the trace to this dynamic, we obtain
$\operatorname{tr}(\dot{Q}(t))=$ $~\operatorname{tr}\left(\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]Q(t)\right)+\operatorname{tr}\left(Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}\right)$
$-\operatorname{tr}\left(Q(t)\Big[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]Q(t)\right)+\operatorname{tr}(a(t,M(t))).$
Since $\nabla_{y}f$, $\nabla_{y}h$, $(\ell\ell^{\top})^{-1}$ and $g$ are bounded, and using the trace inequality (72), we have
$\operatorname{tr}\left(Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))\right)$ $\leq\left|\operatorname{tr}\left(Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))\right)\right|$ (53)
$\leq\left|\left[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\right]^{\top}(t,M(t))\right|\operatorname{tr}\left(Q(t)\right).$ (54)
In addition, the ellipticity of $(\ell\ell^{*})^{-1}$ implies that
$\operatorname{tr}\left(Q(t)\left(\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\right)(t,M(t))Q(t)\right)\geq\operatorname{tr}\left(CQ^{2}(t)\right)\geq\dfrac{C}{n}\operatorname{tr}\left(Q(t)\right)^{2},$ (55)
with the last inequality coming from the property (73). Thus, we can write
$\operatorname{tr}\left(Q(t)\left(\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\right)(t,M(t))Q(t)\right)\leq\dfrac{C}{n}\operatorname{tr}\left(Q(t)\right)^{2}.$ (56)
Finally, $a(t,M(t))$ being bounded, we can find three positive constants $C_{1},~C_{2},~C_{3}$ such that
$\operatorname{tr}(\dot{Q}(t))\leq-C_{1}\operatorname{tr}\left(Q(t)\right)^{2}+C_{2}\operatorname{tr}\left(Q(t)\right)+C_{3}.$ (57)
Denoting $\operatorname{tr}\left(Q(t)\right)$ by $p(t)$, the last inequality reads
$\dot{p}(t)\leq-C_{1}p^{2}(t)+C_{2}p(t)+C_{3}.$ (58)
Then, since $Q(t)$ is bounded, there exists $C_{4}$ such that $C_{2}p(t)+C_{3}\leq C_{4}$. Therefore,
$\dot{p}(t)\leq-C_{1}p^{2}(t)+C_{4}.$ (59)
Then, we can say that the solution of (58) with equality sign will be less than the solution of (59) with the equality sign and both starting at the same initial condition $p(0)$.
Similarly, if $\overline{p}(t)$ denotes the trace of $Q^{-1}(t)$, writing the equation of the dynamic of $Q^{-1}$, and relying on the assumptions that $\nabla_{y}h$ is bounded, $a$ and $(\ell\ell^{\top})^{-1}$ are elliptic, we can find two constants $C^{\prime}_{1}$ and $C^{\prime}_{2}$ such that the trace of $Q^{-1}(t)$ is less than the solution of
$\dot{\overline{p}}(t)=-C^{\prime}_{1}\overline{p}^{2}(t)+C^{\prime}_{2}.$ (60)
To obtain equation (60), we first derive the dynamics of $Q^{-1}$ by taking advantage of the fact that $Q(t)Q^{-1}(t)=I$. This implies that $\dot{Q}(t)Q^{-1}(t)+Q(t)\dot{Q}^{-1}(t)=0$ and thus $\dot{Q}^{-1}(t)=-\left(Q^{-1}\dot{Q}Q^{-1}\right)(t)$.
Hence, substituting $\dot{Q}(t)$ it yields
$\dot{Q}^{-1}(t)=$ $~-Q^{-1}(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))-\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))Q^{-1}(t)$
$~+\Big[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))-Q^{-1}(t)a(t,M(t))Q^{-1}(t).$
Applying the trace and using the same arguments as above, (60) holds.
Now, we have $\operatorname{tr}(Q(t))\leq p(t)$ and $\operatorname{tr}(Q^{-1}(t))\leq\overline{p}(t)$. This implies that
$\overline{p}^{-1}(t)\leq\operatorname{tr}(Q(t))\leq p(t),$ (61)
and then
$\overline{p}^{-1}(t)I\preceq Q(t)\preceq p(t)I.$ (62)
So we only have to prove that $\overline{p}^{-1}(t)\geq Cp(t)$ or equivalently that $p(t)\overline{p}(t)$ is bounded. We have
$\dfrac{d}{dt}\left(p(t)\overline{p}(t)\right)~=$ $~\dot{p}(t)\overline{p}(t)+p(t)\dot{\overline{p}}(t)$
$~=$ $~(-C_{1}p^{2}(t)+C_{2})\overline{p}(t)+p(t)(-C^{\prime}_{1}\overline{p}^{2}(t)+C^{\prime}_{2})$
$~=$ $~(-C_{1}p(t)-C^{\prime}_{1}\overline{p}(t))p(t)\overline{p}(t)+C_{2}\overline{p}(t)+C^{\prime}_{2}p(t)$
$\leq$ $~-C_{3}(p(t)+\overline{p}(t))p(t)\overline{p}(t)+C_{4}(\overline{p}(t)+p(t))$
$\leq$ $~C_{3}(p(t)+\overline{p}(t))(\dfrac{C_{4}}{C_{3}}-p(t)\overline{p}(t))$
$\leq$ $~C_{3}(p(t)+\overline{p}(t))(C_{5}-p(t)\overline{p}(t)),$
for some positive constants $C_{3},C_{4},C_{5}$. Hence, the derivative of $p(t)\overline{p}(t)$ is negative as soon as $p(t)\overline{p}(t)$ is greater than $C_{5}$. Moreover, from the assumption on the eigenvalues of $Q(0)$, $p(0)\overline{p}(0)$ is bounded, thus $p(t)\overline{p}(t)$ is bounded as well. Therefore, the conclusion of Lemma 4.3 holds. ∎
5 Impact of the initial error on the EKF estimation error
To implement the EKF procedure, we must assume that the initial value of the hidden state $Y(0)$ is normally distributed with mean $M(0)$ and covariance matrix $Q^{\varepsilon}(0)$. Therefore, initial guesses for $M(0)$ and $Q^{\varepsilon}(0)$ are required. These initial guesses generate an initial error, which could impact the convergence of the extended Kalman filter to the true signal. It is necessary to analyze how the initial error evolves over time. The theorem below demonstrates that under certain technical conditions, the initial error diminishes exponentially fast.
Theorem 5.1
Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be an extended Kalman filter for the filtering problem (1), with the gain $G$ defined as
$G(t)=\left[g(t)\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\right](\ell\ell^{\top})^{-1}(t,M(t)),$ (63)
where $Q=\frac{1}{\varepsilon}Q^{\varepsilon}$. If the following assumptions are fulfilled,
  1. (A1)
    $\sigma(t,Y(t))$ and $g(t,Y(t))$ are bounded in $L^{\infty-},$
  2. (A2)
    functions $f$ and $h$ are $\pazocal{C}^{1}$ with bounded derivatives,
  3. (A3)
    the observable process $G(t)$ is such that for any $\mathbb{F}^{Z}$-adapted process $\xi$, the process
    $A(t)=\nabla_{y}f(t,\xi(t))-G(t)\nabla_{y}h(t,\xi(t))$ (64)
    is exponentially stable. More precisely, it satisfies (36) for some $c>0$.
    Fixing some $c_{0}<c$, then it holds
    $|Y(t)-M(t)|_{q}\leq C_{q}\sqrt{\varepsilon}+C_{q}|Y(0)-M(0)|_{q}e^{-c_{0}t}.$ (65)
Remark 5.2
Assumption (A1) and (A2) of this theorem are satisfied for the $SI^{\pm}S$ epidemic model, since we consider constant population sizes and strictly positive compartments sizes. For the third assumption, the expression of $A$ is as follows
$A(\tau)=$ $\varepsilon\beta\left[1-2(X_{1}(0)+\varepsilon\xi(\tau))-(X_{2}(0)+\varepsilon Z(\tau))\right]-(\alpha+\rho^{-})\varepsilon$
$-\alpha^{2}\varepsilon\left(Q(\tau)-(X_{1}(0)+\varepsilon\xi(\tau))\right)\left(\alpha(X_{1}(0)+\varepsilon\xi(\tau))+\rho^{-}(X_{2}(0)+\varepsilon Z(\tau))\right)^{-1}.$
Since $A(\tau)$ is a scalar, it is exponentially stable if it is negative for any $\mathbb{F}^{Z}$-adapted process $\xi$. Hence, the parameters and initial values of the filter must be such that $A(\tau)$ is negative in the estimation time window.
The conclusion of Theorem 5.1 shows that for large $t$, the first term $C_{q}\sqrt{\varepsilon}$ is the dominating part of the upper bound on the right-hand side. This implies that for $t$ large enough the filtering error is uniformly bounded w.r.t. time. The proof of Theorem 5.1 involves the use of the following lemma, which is proven in picard1991efficiency.
Lemma 5.3
Consider a family $A(t)$ of locally bounded processes with values in the class of square matrices of order $n$,
  1. 1.
    Suppose that there exist a uniformly bounded and elliptic family $\pazocal{K}(t)$, and a constant number $k>0$ such that $A(t)$ is $(\pazocal{K}(t),k)$ stable, then $A(t)$ is exponentially stable and the estimate (36) is satisfied with $c=k/2$.
  2. 2.
    Conversely, if $A(t)$ is uniformly bounded and exponentially stable (so that estimate (36) holds for some $c>0$) then for any $k<2c$, there exists a family of uniformly bounded and elliptic processes $\pazocal{K}(t)$ that are adapted to the filtration generated by $A$ and are such that $A(t)$ is $(\pazocal{K}(t),k)$ stable. More generally, if $q_{0}$ is a family of uniformly positive numbers, there exists a family of uniformly elliptic processes $\pazocal{K}(t)$ such that $\pazocal{K}(0)=q_{0}I,$ $A(t)$ is $(\pazocal{K}(t),k)$ stable and
    $\operatorname{tr}(\pazocal{K}(t))\leq C\left(1+q_{0}e^{(-(2c-k)t)}\right).$ (66)
Proof
(of Theorem 5.1) Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be the extended Kalman filter of our model with gain $G(t)$ defined in the theorem and initial guess $M(0)$. Since $f$ and $h$ are $C^{1}$ with bounded derivatives, there exists a $\mathbb{F}$-adapted process $\xi$ such that
$f(t,Y(t))-f(t,M(t))=\nabla_{y}f(t,\xi(t))\left(Y(t)-M(t)\right),$ (67)
and
$h(t,Y(t))-h(t,M(t))=\nabla_{y}h(t,\xi(t))\left(Y(t)-M(t)\right),$ (68)
Thus, defining $\pazocal{X}(t):=Y(t)-M(t)$, it follows that
$f(t,Y(t))-f(t,M(t))-G(t)\left(h(t,Y(t))-h(t,M(t))\right)=$ $\left(\nabla_{y}f(t,\xi(t))-G(t)\nabla_{y}h(t,\xi(t))\right)\pazocal{X}(t)$
$=$ $A(t)\pazocal{X}(t),$
with $A(t):=\nabla_{y}f(t,\xi(t))-G(t)\nabla_{y}h(t,\xi(t)).$
Therefore, defining $\widetilde{W}(t):=\left(W^{(1)}(t),W^{(2)}(t)\right)^{\top}$, we have
$d\pazocal{X}(t)$ $=d\left(Y(t)-M(t)\right)$
$=\left[f(t,Y(t))-f(t,M(t))-G(t)\left(h(t,Y(t))-h(t,M(t))\right)\right]dt+$
$\phantom{=}~~\sqrt{\varepsilon}\left[\sigma(t,Y(t))dW^{(1)}(t)+\left(g(t,Y(t))-G(t)\ell(t,Y(t))\right)dW^{(2)}(t)\right]$
$\phantom{=}~~\left(\nabla_{y}f(t,\xi_{t})-G(t)\nabla_{y}h(t,\xi(t))\right)\pazocal{X}(t)dt+$
$\phantom{=}~~\sqrt{\varepsilon}\left(\sigma(t,Y(t)),~g(t,Y(t))-G(t)\ell(t,Y(t))\right)d\widetilde{W}(t)$
$=A(t)\pazocal{X}(t)dt+\sqrt{\varepsilon}D(t)d\widetilde{W}(t),$
with $D(t):=\left(\sigma(t,Y(t)),~g(t,Y(t))-G(t)\ell(t,Y(t))\right).$
Since $A(t)$ is exponentially stable and bounded by assumption, Lemma 5.3 asserts that for any $k<2c$, there exists a family of processes $\pazocal{K}(t)$ that are uniformly bounded, elliptic and adapted such that $A(t)$ is $(\pazocal{K}(t),k)$ stable, and for some family of uniformly positive numbers $q_{0}=|Y(0)-M(0)|_{q}^{2}$, $\pazocal{K}(t)$ satisfy
$\operatorname{tr}(\pazocal{K}(t))\leq C\left(1+q_{0}e^{(-(2c-k)t)}\right).$ (69)
We deduce also that $A(t)$ is $(\varepsilon\pazocal{K}(t),k)$ stable. We are then in the setting of Lemma 4.4 with $F(t)=0$. Thus for $\alpha=0$ and because $\sigma,~~g$ and $\ell$ are bounded, for any $q\geq 1$ and $\varepsilon$ small enough, if $\pazocal{X}^{\top}(0)(\varepsilon\pazocal{K}(0))^{-1}\pazocal{X}(0)$ is bounded in $L^{q}$, then the process $\pazocal{X}^{\top}(t)(\varepsilon\pazocal{K}(t))^{-1}\pazocal{X}(t)$ is bounded in $L^{q}$. In addition,
$\varepsilon\pazocal{K}(t)\preceq\operatorname{tr}(\varepsilon\pazocal{K}(t))I$~~ $\text{implies}~~(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}I\preceq(\varepsilon\pazocal{K}(t))^{-1},$
$\text{thus}~~~\pazocal{X}^{\top}(t)(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}\pazocal{X}(t)\leq\pazocal{X}^{\top}(t)(\varepsilon\pazocal{K}(t))^{-1}\pazocal{X}(t),$
$\text{then}~~~\pazocal{X}^{\top}(t)(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}\pazocal{X}(t)\text{ is bounded in }L^{q}.$
And finally, given the fact that $\pazocal{X}^{\top}(t)(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}\pazocal{X}(t)=|(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1/2}\pazocal{X}(t)|^{2}$, we deduce that $(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1/2}\pazocal{X}(t)$ is bounded in $L^{q}$.
Therefore, there exists $C_{q}$ such that
$|(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1/2}\pazocal{X}(t)|_{q}\leq C_{q},$ (70)
This implies, using the inequality $(1+x)^{1/2}\leq 1+x^{1/2}$ for $x\geq 0$, that
$|\pazocal{X}(t)|_{q}$$\leq$ $\sqrt{\varepsilon}C_{q}(\operatorname{tr}(\pazocal{K}(t)))^{1/2}$
$\leq$ $\sqrt{\varepsilon}C_{q}\left(1+q_{0}e^{(-(2c-k)t)}\right)^{1/2}$
$\leq$ $C_{q}\sqrt{\varepsilon}+C_{q}\sqrt{\varepsilon}|Y(0)-M(0)|_{q}e^{(-(c-k/2)t)}$
$\leq$ $C_{q}\sqrt{\varepsilon}+C_{q}|Y(0)-M(0)|_{q}e^{(-c_{0}t)},$
with $c_{0}=c-k/2$ and $\varepsilon$ small enough. ∎
Acknowledgements.
The authors thank the collaborators within the DFG research project “CESMO - Contain Epidemics with Stochastic Mixed-Integer Optimal Control”, in particular Gerd Wachsmuth, Armin Fügenschuh, Markus Friedemann, Jesse Beisegel (BTU Cottbus–Senftenberg), for insightful discussions and valuable suggestions that improved this paper. Funding  The authors gratefully acknowledge the support by the Deutsche Forschungsgemeinschaft (DFG), award number 458468407, and by the German Academic Exchange Service (DAAD), award number 57417894.
Appendix
Appendix A The Moore-Penrose pseudoinverse
Let $n,m$ be two positive integers, and $\pazocal{M}_{n,m}(\mathbb{R})$ the class of $n\times m$ matrices with real entries. Every $A\in\pazocal{M}_{n,m}(\mathbb{R})$ has a pseudoinverse and, moreover, the pseudoinverse, denoted $A^{+}\in\pazocal{M}_{n,m}(\mathbb{R})$, is unique. A purely algebraic characterization of $A^{+}$ is given in the next theorem proved by Penrose in 1956.
Theorem A.1
Let $A\in\pazocal{M}_{n,m}(\mathbb{R})$. Then $B=A^{+}$ if and only if
  1. 1.
    $ABA=A$,
  2. 2.
    $BAB=B$,
  3. 3.
    $(AB)^{\top}=AB$,
  4. 4.
    $(BA)^{\top}=BA$.
Furthermore, $A^{+}$ always exists and is unique
Appendix B General results on symmetric positive definite (semi-definite) matrices
We refer to the book horn2012matrix for further details and proof of results in this section. Let $n,m$ be two positive integers, and $\pazocal{M}_{n}(\mathbb{R})$ the class of $n\times n$ matrices with real elements:
Definition B.1
A symmetric matrix $A\in\pazocal{M}_{n}(\mathbb{R})$ is
  1. 1.
    positive definite if $x^{\top}Ax>0$ for all nonzero $x\in\mathbb{R}^{n}$;
  2. 2.
    positive semi-definite if $x^{\top}Ax\geq 0$ for all nonzero $x\in\mathbb{R}^{n}$.
Proposition B.2
Let $A_{1},\ldots,A_{k}\in\pazocal{M}_{n}(\mathbb{R})$ be positive semi-definite and let $\alpha_{1},\ldots,\alpha_{k}$ be nonnegative real numbers. Then $\sum_{i=1}^{k}\alpha_{i}A_{i}$ is positive semi-definite. If there is a $j\in\{1,\ldots,k\}$ such that $\alpha_{j}>0$ and $A_{j}$ is positive definite, then $\sum_{i=1}^{k}\alpha_{i}A_{i}$ is positive definite.
Proposition B.3
Let $A\in\pazocal{M}_{n}(\mathbb{R})$ be positive semi-definite (respectively, positive definite). Then $\operatorname{tr}A$, $\operatorname{det}A$, and the principal minors of $A$ are all nonnegative (respectively, positive). Moreover, $\operatorname{tr}A=0$ if and only if $A=0$.
Proposition B.4
Let $A\in\pazocal{M}_{n}(\mathbb{R})$ be positive semi-definite and let $x\in\mathbb{R}^{n}$. Then $x^{\top}Ax=0$ if and only if $Ax=0$.
Proposition B.5
A positive semi-definite matrix is positive definite if and only if it is nonsingular.
Theorem B.6
Let $A\in\pazocal{M}_{n}(\mathbb{R})$ be symmetric. Then $x^{\top}Ax>0$ (respectively, $x^{\top}Ax\geq 0$) for all nonzero $x\in\mathbb{R}^{n}$ if and only if every eigenvalue of $A$ is positive (respectively, nonnegative).
Theorem B.7
A symmetric matrix is positive semi-definite if and only if all of its eigenvalues are nonnegative. It is positive definite if and only if all of its eigenvalues are positive.
Corollary B.8
A nonsingular symmetric matrix $A\in\pazocal{M}_{n}(\mathbb{R})$ is positive definite if and only if $A^{-1}$ is positive definite.
Corollary B.9
If $A\in\pazocal{M}_{n}(\mathbb{R})$ is positive semi-definite, then so is each $A^{k},~k=1,2,\ldots$.
Theorem B.10
Let $A\in\pazocal{M}_{n}(\mathbb{R})$ be symmetric.
  1. 1.
    A is positive semi-definite if and only if there is a $B\in\pazocal{M}_{m,n}(\mathbb{R})$ such that $A=B^{\top}B$.
  2. 2.
    If $A=B^{\top}B$ with $B\in\pazocal{M}_{m,n}$, and if $x\in\mathbb{R}^{n}$, then $Ax=0$ if and only if $Bx=0$, so $\operatorname{nullspace}A=\operatorname{nullspace}B$ and $\operatorname{rank}A=\operatorname{rank}B$.
  3. 3.
    If $A=B^{\top}B$ with $B\in\pazocal{M}_{m,n}(\mathbb{R})$, then $A$ is positive definite if and only if $B$ has full column rank.
Appendix C The Loewner partial order
Proposition C.1
If $A\in\pazocal{M}_{n}(\mathbb{R})$ is symmetric with the smallest and largest eigenvalues $\lambda_{min}(A)$ and $\lambda_{max}(A)$, respectively, then $\lambda_{max}(A)I\succeq A\succeq\lambda_{min}(A)$
Proposition C.2
Let $A\in\pazocal{M}_{n}(\mathbb{R})$ be symmetric,
  1. 1.
    $I\succeq A$ if and only if $\lambda_{max}(A)\leq 1$;
  2. 2.
    $I\succ A$, if and only if $\lambda_{max}(A)<1$.
Definition C.3
Let $X\in\pazocal{M}_{n,m}(\mathbb{R})$. Then $\sigma_{1}(X)=\lambda_{max}^{1/2}(XX^{\top})=\lambda_{max}^{1/2}(X^{\top}X)=\sigma_{1}(X^{\top})$ is the largest singular value (the spectral norm) of $X$. We say that $X$ is a contraction if $\sigma_{1}(X)\leq 1$; it is a strict contraction if $\sigma_{1}(X)<1$.
Theorem C.4
Let $A,~B\in\pazocal{M}_{n}(\mathbb{R})$ be symmetric and let $S\in M_{n,m}$. Then
  1. 1.
    if $A\succeq B$, then $S^{\top}AS\succeq S^{\top}BS$;
  2. 2.
    if $\operatorname{rank}S=m$, then $A\succ B$ implies $S^{\top}AS\succ S^{\top}BS$
  3. 3.
    if $m=n$ and $S\in M_{n}$ is nonsingular, then $A\succ B$ if and only if $S^{\top}AS\succ S^{\top}BS$; $A\succeq B$ if and only if $S^{\top}AS\succeq S^{\top}BS$;
  4. 4.
    $I_{m}\succ S^{\top}S$ (respectively, $I_{n}\succ SS^{\top}$) if and only if $S$ is a strict contraction; $I_{m}\succeq S^{\top}S$ (respectively, $I_{n}\succeq SS^{\top}$) if and only if $S$ is a contraction;
Theorem C.5
Let $A,B\in\pazocal{M}_{n}(\mathbb{R})$ be symmetric. Let $\lambda_{1}(A)\leq\ldots\leq\lambda_{n}(A)$ and $\lambda_{1}(B)\leq\ldots\leq\lambda_{n}(B)$ be the ordered eigenvalues of A and B, respectively.
  1. 1.
    If $A\succ 0$ and $B\succ 0$, then $A\succeq B$ if and only if $B^{-1}\succeq A^{-1}$.
  2. 2.
    If $A\succ 0$, $B\succeq 0$, and $A\succeq B$ , then $A^{1/2}\succeq B^{1/2}$.
  3. 3.
    If $A\succeq B$, then $\lambda_{i}(A)\geq\lambda_{i}(B)$ for each $i=1,\ldots,n$.
  4. 4.
    If $A\succeq B$, then $\operatorname{tr}A\geq\operatorname{tr}B$ with equality if and only if $A=B$.
  5. 5.
    If $A\succeq B\succeq 0$, then $\operatorname{det}A\geq\operatorname{det}B\geq 0$.
Theorem C.6
Let $A\in\pazocal{M}_{n}(\mathbb{R})$ be positive semi-definite and let $B\in\pazocal{M}_{n}(\mathbb{R})$ be symmetric. The following four statements are equivalent:
  1. 1.
    $x^{\top}Ax\geq|x^{\top}Bx|$ for all $x\in\mathbb{C}^{n}$.
  2. 2.
    $x^{\top}Ax+y^{\top}Ay\geq 2|x^{\top}By|$ for all $x,y\in\mathbb{C}^{n}$.
  3. 3.
    $H=\left[\begin{array}[]{cc}A&B\\ B&A\end{array}\right]$ is positive semi-definite.
  4. 4.
    There is a symmetric contraction $X\in\pazocal{M}_{n}(\mathbb{R})$ such that $B=A^{1/2}XA^{1/2}$.
    If A is positive definite, then the following statement is equivalent to 1:
  5. 5.
    $\rho(A^{-1}B)\leq 1$.
Appendix D Some properties of the trace
We refer to gabih2020asymptotic for the proof of the following lemma.
Lemma D.1
(Properties of symmetric and positive semi-definite matrices) Let $A,B\in\pazocal{M}_{d}(\mathbb{R})$, $d\in\mathbb{N}$, symmetric and positive semi-definite matrices. Then it holds
  1. 1.
    $\lambda_{min}(A)\operatorname{tr}(B)\leq\operatorname{tr}(AB)\leq\lambda_{max}(A)\operatorname{tr}(B)$ (71)
    where $\lambda_{min}(A)$ and $\lambda_{max}(A)$ denote the smallest and largest eigenvalue of A, respectively.
  2. 2.
    $\dfrac{\operatorname{tr}(B)}{\operatorname{tr}\left(A^{-1}\right)}\leq\operatorname{tr}(AB)\leq\operatorname{tr}(A)\operatorname{tr}(B)$ (72)
    where for the first inequality A is assumed to be positive definite.
  3. 3.
    $\operatorname{tr}^{2}(A)\geq\operatorname{tr}\left(A^{2}\right)\geq\dfrac{1}{d}\operatorname{tr}^{2}(A)$ (73)
  4. 4.
    $|A|_{F}=\sqrt{\operatorname{tr}\left(A^{2}\right)}\leq\operatorname{tr}(A)$ (74)
    where $|A|_{F}$ denotes the Frobenius norm of A.
Theorem D.2
Let $A,B\in\pazocal{M}_{d}(\mathbb{R})$, $d\in\mathbb{N}$, symmetric and positive semi-definite matrices, then
$\operatorname{tr}(AB)\leq|\operatorname{tr}(AB)|\leq|A|_{2}\operatorname{tr}(B)$ (75)
where $|A|_{2}$ denotes the spectral norm or largest singular value of $A$.